Kimberly Fields has $25,000 invested in two accounts, part at 8.5% simple interest and part at 4% simple interest for a period of 1 year. How much did she invest at each rate if each account earned the same interest?
Answer by richwmiller(17219)

We know the total amount of money invested. $25000
x+y=25000,
We know that the difference in interest earned by the two accounts is $0
0.085*x-0.04*y=0
x=25000-y
We substitute for x
0.085*(25000-y)-0.04*y=0
We multiply out
2125.0-0.085y-0.04*y=0
We combine like terms.
2125.0=0.125*y
Isolate y
y=2125.0/0.125
y=17000 at 4%
Calculate x
x=25000-17000
x=8000 at 8.5%
Interest earned at 8.5% is 680.0
Interest earned at 4% is 680
We check
0.085*8000-0.04*17000=0
680.0-680=0
0.0=0
Since this statement is TRUE and neither x nor y is negative then all is well
codeintdy